OpenStudy (anonymous):

If 3.0 L of oxygen gas at 177°C is cooled at constant pressure until the volume becomes 1.50 L, then what is the final temperature

OpenStudy (anonymous):

Volume is directly proportional to kelvin temp

OpenStudy (anonymous):

so if volume is halved...

OpenStudy (anonymous):

i always sudgest to start from what you know and move forward so you dont have to remember equations... so lets start with pV=nRT now since there is no change in amount you can write next assumption: n1=n2 - where n1 is state before change and n2 state after change now you can write next equation: p1V1/RT1 = p2V2/RT2 since you have Rydbergs constant on both sides they can be crossed out and you get: p1V1/T1 = p2V2/T2 now you have in task that pressure is constant so it crosses out like Rydbergs constant and you get: V1/T1 = V2/T2 now all you need to do is little math and rearange the equation and you get: T2=V2T1/V1 and that is it... now you just put in the numbers

OpenStudy (anonymous):

You might remember Charles' Law which states that the volume of a fixed mass of gas is directly proportional to its Kelvin temperature or \[V _{1}/T _{1}=V _{2}/T _{2}\] Remember to change the temperature to Kelvin. Volume can be in any units as long as both volumes are in the same units.